SPECTRAL INHOMOGENEITY AS A DISCRETE, COUNTABLE ENTITY – THE ROLE OF RESIDUAL ORTHOHYDROGEN MOLECULES ON THE ROVIBRATIONAL SPECTRA OF DOPANTS IN PARAHY-DROGEN SOLIDS

\underline{MARIO E. FAJARDO} and SIMON TAM, AFRL/PRSP, Bldg. 8451, Propulsion Sciences and Advanced Concepts Division, US Air Force Research Lab, Edwards AFB, CA 93524-7680..

We report high resolution (0.005 cm^{−1}) infrared (IR) absorption spectra of doped millimeters-thick parahydrogen (pH_{2}) solids with intentionally elevated residual orthohydrogen (oH_{2}) concentrations. In addition to the weakly-hindered rotor spectrum of isolated H_{2}O molecules, H_{2}O/oH_{2}/pH_{2} samples exhibit a series of sharp (0.03 cm^{−1} FWHM) absorption peaks in the 1593 to 1597 cm^{−1} region which we show are due to H_{2}O-oH_{2} clusters. We have also identified CO_{2}-oH_{2} clusters as the source of the very sharp (0.007 cm^{−1} FWHM) fine structure in IR spectra of CO_{2}/oH_{2}/pH_{2} solids.^{a} The gradual evolution of the CO_{2}-oH_{2} features with increasing oH_{2} concentration raises the interesting notion of spectral inhomogeneity as a discrete, countable entity.

aS. Tam and M.E. Fajardo, Low Temp. Phys. v26, p653 (1998).